OpenStudy (joshoyen):

Write the slope-intercept form of the equation of each line. 1. Through: (-1,-1) and (5,2)

pooja195 (pooja195):

Start by finding the slope use the slope formula \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (joshoyen):

3/6

OpenStudy (joshoyen):

= 1/2

pooja195 (pooja195):

Good now now we use the slope intercept form y=mx+b We know that m=slope y=1/2x+b take one of the coordinates given and solve for b

OpenStudy (joshoyen):

hm, im confused on this part

pooja195 (pooja195):

Ok so pick a coordinate (given in the problem)

OpenStudy (joshoyen):

(-1,-1)

pooja195 (pooja195):

Ok we take those coordinates and plug them into the equation -1=1/2(-1)+b solve for b

OpenStudy (joshoyen):

Ohhh okay

OpenStudy (joshoyen):

thanks!

pooja195 (pooja195):

Can you please solve it :) ?

OpenStudy (joshoyen):

wait I got -1=-1/2+b

pooja195 (pooja195):

Add 1/2 to both sides

OpenStudy (joshoyen):

-1/2=b ?

pooja195 (pooja195):

Good so you have y=1/2x-1/2 :P

OpenStudy (joshoyen):

Ohh okay

OpenStudy (joshoyen):

now i get it, thanks!

pooja195 (pooja195):

yw ^_^

OpenStudy (joshoyen):

Could you also help me on this problem. Through: (-1,2), slope = -6

OpenStudy (joshoyen):

@pooja195

pooja195 (pooja195):

@jordynhazan

pooja195 (pooja195):

Still need help?

OpenStudy (joshoyen):

Oh i'm good thanks , sorry for the late reply @pooja195

pooja195 (pooja195):

Its alright :-)